OpenStudy (anonymous):

what is the equation of the line, in slope-intercept form, that goes through the point (-6,_4) with m=-1/2? A) y= -1/2x+7 B) y= -1/2x+1 C) y= -1/2x-1 D)y= -1/2x-7

OpenStudy (mertsj):

Plug into point-slope form which is: \[y-y _{1}=m(x-x _{1})\] where \[x _{1}=-6, y _{1}=-4, m = \frac{-1}{2}\]

OpenStudy (anonymous):

would it be C?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

thank you!

OpenStudy (mertsj):

Wait. I might have messed up a sign. Let me check.

OpenStudy (anonymous):

ok

OpenStudy (mertsj):